A horizontal wire of length 1.0 m carries a curren
Answer #1

length of horizontal wire = 1m ; current ( left to right ) =48A

second wire length =0.5m ; mass=0.5Kg ; height from 1st wire =0.30 m

(a) direction of current in second suspended wire = From Right to left ( because conductors carrying current in opposite directions repel each other)

(b) field at upper wire due to bottom wire

   B = \mu _{0}i/2\pi r    i = current in first wire r = distance of second wire from first

   \mu _{0} = permeability constant = 1.26 X 10^(-6) henry/m

B = 1.26 X 10^(-6) H/m X 48 A / 2 X 3.14 X 0.3 m = 32.1 \mu T

(c) weight of the top wire that must be balanced = m g = 0.5 Kg X 9.8m/s^2 = 4.9 N

(d) Current in top wire to balance weight = ?

F = 4.9 N ( to balance wight of top wire) is required to be created at top wire due to bottom wire. This should be repulsive force

F = \mu _{0}l i_{a}i_{b}/2\pi d  where l = length of 2nd wire at top

i_{a} = current in wire at bottom , i_{b} = current in top wire d= distance between the wires

F = 1.26 X 10^(-6) H/m X 0.5 m X 48 A X i_{b} /(2\pi \times 0.3m) = 4.9N

i_{b} = 2 X 3.14 X 0.3 m X 49 N /( 1.26 X 10^(-6) X ).5 X 48 A) = 3.0 X 10^(6) A